# Lagrangian chaos and unique ergodicity for stochastic primitive equations

### v1

@@ 6. Strong Feller for the Lagrangian and projective processes
This section is organized as follows. In Subsection 6.1, we reduce the proof of the strong Feller property for the original system to a truncated one. The latter is then further reduced to a suitable gradient estimate for the truncated problem. Subsection 6.2 introduces the necessary notation for Malliavin calculus. Subsection 6.3 provides an elementary proof of Proposition 5.6, which also sets the stage for the more intricate proof of Proposition 3.11. Finally, in Subsections 6.4 and 6.5, we prove the required gradient estimate for the projective process $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ and the estimate for the inverse of the reduced Malliavin matrix, respectively.
@@ 6.1. Strong Feller of the projective process via its cutoff
where $W^{(v)}_{t}=W_{t}$ is the cylindrical Brownian motion on $\mathbb{L}^{2}_{0}$, described in Subsection 2.2.2, $W_{t}^{(\eta)}$ is a ${\mathbb{R}}^{6}$-Brownian motion in ${\mathbb{R}}^{6}$ independent of $W^{(v)}_{t}$ and $Q$ is as Subsection 2.2.2. Moreover, as above, $u=(v,w)$ where $w$ is as in (2.1). The system (6.1) is complemented with the initial conditions $v_{0}=v\in\mathbb{H}^{\sigma}_{0}$, $\mathbf{x}_{0}=\mathbf{x}\in\mathbb{T}^{3}$, $\xi_{0}=\xi\in\mathbb{S}^{2}$ and $\eta_{0}=\eta\in{\mathbb{R}}^{6}$.
In contrast to Subsection 3.3, it will be more convenient here to treat $(\xi_{t})_{t}$ as an $\mathbb{S}^{2}$-valued process. However, to establish the strong Feller property as in Proposition 3.11, this approach is sufficient. Indeed, any observable $\psi:\mathbb{H}^{\sigma}\times\mathbb{T}^{3}\times\mathrm{Pr}_{2}\to{\mathbb{R}}$ can be lifted to an obsevable $\widetilde{\psi}$ on $\mathbb{H}^{\sigma}\times\mathbb{T}^{3}\times\mathbb{S}^{2}$ via the formula $\widetilde{\psi}(v,\mathbf{x},\xi)=\psi(v,\mathbf{x},\pi(\xi))$, where $\pi:\mathbb{S}^{2}\to\mathrm{Pr}_{2}$ is the standard quotient map.
@@
$V_{t}\stackrel{{\scriptstyle{\rm def}}}{{=}}\begin{pmatrix}v_{t},\mathbf{x}_{t},\xi_{t},\eta_{t}\end{pmatrix}^{\top}\in\mathcal{H}\qquad\text{ where }\qquad\mathcal{H}\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbb{H}^{\sigma}_{0}\times\mathbb{T}^{3}\times\mathbb{S}^{2}\times{\mathbb{R}}^{6}.$
@@
(6.3) $\chi(t)=\left\{\begin{aligned} &0&\quad&\text{ for }t\leq 1,\\ &1&\quad&\text{ for }t\geq 2.\end{aligned}\right.$
@@
(6.4) $F^{(r)}(V)\stackrel{{\scriptstyle{\rm def}}}{{=}}\big{[}1-\chi_{3r}(\|v\|_{H^{\sigma}})\big{]}F(V)+\chi_{r}(\|v\|_{H^{\sigma}})H(\xi,\eta),$
where $V=(v,\mathbf{x},\xi,\eta)$ and
$H(\xi,\eta)=\begin{pmatrix}0\\ \sum_{j=1}^{3}a_{j}\frac{\eta_{j}}{\sqrt{1+|\eta_{j}|^{2}}}\\ \Pi_{\xi}\sum_{j=1}^{3}a_{j}\frac{\eta_{j+3}}{\sqrt{1+|\eta_{j+3}|^{2}}}\\ 0\end{pmatrix}$
and $(a_{j})_{j=1}^{3}$ is the canonical basis of ${\mathbb{R}}^{3}$. The idea of the above truncation is that some ‘randomness’ is transferred from the $\eta$-equation to $\mathbf{x}_{t}$ and $\xi_{t}$ via the nonlinearity $\chi_{r}(\|v\|_{H^{\sigma}})H(\xi,\eta)$ when $\|v\|_{\mathbb{H}^{\sigma}}$ is large. However, for small values of $\|v\|_{\mathbb{H}^{\sigma}}$, the coupling is not efficient enough, and a hypoellipticity argument has to be exploited; see Subsection 6.5.1.
@@
where $V\in\mathcal{H}$. The global well-posedness of the stochastic PEs with cutoff (6.5) follows from standard arguments due to the presence of the truncation and $\sigma\geq 2$. For all $r,t\geq 0$ and bounded Borel maps $\phi:\mathcal{H}\to{\mathbb{R}}$, we define the semigroups
@@
where $V_{t}$ and $V_{t}^{(r)}$ is the solution to (6.2) and (6.5) with initial data $V$, respectively. We say that the process $(V_{t})_{t}$ (resp. $(V_{t}^{(r)})_{t}$) is strong Feller if $P_{t}\phi\in C(\mathcal{H};{\mathbb{R}})$ (resp. $P_{t}^{(r)}\phi\in C(\mathcal{H};{\mathbb{R}})$) for all bounded Borel maps $\phi:\mathcal{H}\to{\mathbb{R}}$.
The next result allows us to deduce the strong Feller property of $(V_{t})_{t\geq 0}$ to the one of $(V^{(r)}_{t})_{t\geq 0}$ for sufficiently large $r$.
@@
Fix a bounded Borel observable $\phi:\mathcal{H}\to{\mathbb{R}}$, a time $t\geq 0$ and an initial data $V\in\mathcal{H}$. To prove the claim, we need to show that, for each $\varepsilon_{0}>0$, there exists $\delta_{0}=\delta_{0}(\varepsilon_{0})>0$ such that
@@
Let $V=(v,\mathbf{x},\xi,\eta)\in\mathbb{H}^{\sigma}_{0}\times M$. By (6.1), it is clear that $V_{t}=(v_{t},\mathbf{x}_{t},\xi_{t},\eta+W^{(\eta)}_{t})$ for $t\geq 0$, where $v_{t}$ is the unique solution to (1.1) and $(\mathbf{x}_{t},\xi_{t})$ the corresponding processes uniquely determined by $v_{t}$ as in (6.1). By Theorem 5.1(1), we can choose $r_{1}\geq r_{0}$ such that, for $\widehat{v}\in\mathbb{H}^{\sigma}_{0}$ satisfying $\|v-\widehat{v}\|_{\mathbb{H}^{\sigma}_{0}}\leq 1$, the solution $\widehat{v}_{t}$ of the stochastic PEs (1.1) with initial data $\widehat{v}\in\mathbb{H}^{\sigma}_{0}$ satisfies
$\mathbf{P}\Big{(}\sup_{s\in[0,t]}\|\widehat{v}_{s}\|_{\mathbb{H}^{\sigma}}\geq r_{1}\Big{)}\leq\frac{\varepsilon}{4\|\phi\|_{L^{\infty}}}.$
@@
we have $\mathbf{P}(\widehat{\tau}<t)\leq\varepsilon/(4\|\phi\|_{L^{\infty}})$. By uniqueness of solutions to the primitive equations (cf., Definition 2.1), it follows that $v^{(r)}=v$ on $[0,\tau)\times\Omega$. Hence, for all $V=(v,\mathbf{x},\xi,\eta)\in\mathcal{H}$ such that $\|\widehat{v}\|_{\mathbb{H}^{\sigma}}\leq N$ and $r\geq r_{1}$,
@@
Finally, for all $r\geq r_{1}$ and $V^{\prime}\in\mathcal{H}$ satisfying $\|V-V^{\prime}\|_{\mathcal{H}}\leq 1$,
@@
In light of Lemma 6.1, it is enough to prove the Markov process $(V^{(r)}_{t})_{t}$ is strong Feller for $r$ sufficiently large. The latter fact is a consequence of the following result. In this formulation, we use the following notation:
@@
$|DP^{(r)}_{t}\phi(V)h|\lesssim t^{-a_{0}}(1+\|V\|_{\mathcal{H}}^{b_{0}})\|\phi\|_{L^{\infty}}\|h\|_{T_{V}\mathcal{H}}$
@@
The fact that the above implies the strong Feller of $(V^{(r)}_{t})_{t}$ with $r\geq r_{0}$ is standard, see e.g., [25] or [16, Proposition 6.1]. For brevity, we omit the details.
In the rest of this section, we focus on proving Proposition 6.2. Our proof of Proposition 6.2 follows the line of the of [16, Proposition 6.1] done in Section 6 of the same work. However, adjustments are needed for the stochastic PEs (1.1). In particular, because of the noise degeneracy of the third component $w_{t}$ of the vector field $u_{t}$ (see (2.1)), which creates additional difficulties in the study of the invertibility of the partial Malliavin matrix, see Subsection 6.5.1.
@@ 6.2.1. Basics on Malliavin calculus
In this subsection, we introduce the key concepts in the context of Malliavin calculus, primarily to fix the notation for the subsequent subsections. For a more detailed discussion, the reader is referred to the monograph [61], and for extensions to the Banach-valued case, see [65]. Malliavin calculus is typically introduced in the context of isonormal Gaussian processes [61, Definition 1.1.1]. For our purposes, we focus here on $\mathscr{U}$-cylindrical Brownian motion $\mathscr{W}$ on a separable Hilbert space; see e.g., [8, Definition 2.11] for the definition. Recall that such a cylindrical process is equivalent to an isonormal Gaussian process on $L^{2}(0,T;\mathscr{U})$. To see the correspondence between the latter situation and the one analysed here, let us mention that the sequence of standard independent Brownian motions $(W^{(\mathbf{k},\ell)})_{(\mathbf{k},\ell)\in\mathcal{Z}}$ in (2.12) uniquely induced a $\mathbb{L}^{2}_{0}$-cylindrical Brownian motion via the formula: For $f\in\mathbb{L}^{2}_{0}$ and $0\leq s<t\leq T$,
@@
Using the previous identification, the Malliavin derivative $\mathscr{D}$ can be defined as in [61, Subsection 1.2] or [65, Subsection 2.2] on an Hilbert space $\mathscr{H}$ as an unbounded linear operator
@@
In general, computing the Malliavin derivative is not straightforward. Below, we will need to compute the Malliavin derivative of $V_{T}$, where $V_{T}$ is the solution to the truncated SPDE (6.2) for a fixed $r>0$. However, as shown in [61, Subsection 1.2.1], when $\Omega$ is the standard space of paths, the Malliavin derivative of the random variable $V_{T}$ be computed using the Fréchet derivative of the Itô-map (if available) associated to (6.2), i.e., the map which associated to the noise realisation $W$ the solution $V_{T}(W)$, see [61, Subsection 1.2]. To make this concrete, we need to introduce some more notation. To this end, as in Subsection 3.3 and without loss of generality, we may assume that
@@
Using (6.13), the Fréchet differentiability of the Itô map can be demonstrated as in [41, Proposition 4.1]. For the equality in (6.14), the proof can be obtained by arguing as in [61, Subsection 1.2.1].

### v2

@@ 6. Strong Feller for the Lagrangian and projective processes
This section is organised as follows. In Subsection 6.1, we reduce the proof of the strong Feller property for the original system to a truncated one. The latter is then further reduced to a suitable gradient estimate for the truncated problem. Subsection 6.2 introduces the necessary notation for Malliavin calculus. Subsection 6.3 provides an elementary proof of Proposition 5.6, which also sets the stage for the more intricate proof of Proposition 3.11. Finally, in Subsections 6.4 and 6.5, we prove the required gradient estimate for the projective process $(v_{t},\mathbf{x}_{t},\xi_{t})_{t}$ and the estimate for the inverse of the reduced Malliavin matrix, respectively.
@@ 6.1. Strong Feller of the projective process via its cutoff
where $W^{(v)}_{t}=W_{t}$ is the cylindrical Brownian motion on $\mathbb{L}^{2}_{0}$, described in Subsection 2.2.2, $W_{t}^{(\eta)}$ is a ${\mathbb{R}}^{6}$-Brownian motion in ${\mathbb{R}}^{6}$ independent of $W^{(v)}_{t}$ and $Q$ is as Subsection 2.2.2. Moreover, as above, $u=(v,w)$ where $w$ is as in (2.1). The system (6.1) is complemented with the initial conditions $v_{0}=v\in\mathbb{H}^{\sigma}_{0}$, $\mathbf{x}_{0}=\mathbf{x}\in\mathcal{O}$, $\xi_{0}=\xi\in\mathbb{S}^{2}$ and $\eta_{0}=\eta\in{\mathbb{R}}^{6}$.
In contrast to Subsection 3.3, for convenience, here we treat $(\xi_{t})_{t}$ as an $\mathbb{S}^{2}$-valued process. However, to establish the strong Feller property as in Proposition 3.11, this approach is sufficient. Indeed, any observable $\psi:\mathbb{H}^{\sigma}_{0}\times\mathcal{O}\times\mathrm{Pr}_{2}\to{\mathbb{R}}$ can be lifted to an obsevable $\widetilde{\psi}$ on $\mathbb{H}_{0}^{\sigma}\times\mathcal{O}\times\mathbb{S}^{2}$ via the formula $\widetilde{\psi}(v,\mathbf{x},\xi)=\psi(v,\mathbf{x},\pi(\xi))$, where $\pi:\mathbb{S}^{2}\to\mathrm{Pr}_{2}$ is the standard quotient map.
@@
$V_{t}\stackrel{{\scriptstyle{\rm def}}}{{=}}\begin{pmatrix}v_{t},\mathbf{x}_{t},\xi_{t},\eta_{t}\end{pmatrix}^{\top}\in\mathcal{H}\qquad\text{ where }\qquad\mathcal{H}\stackrel{{\scriptstyle{\rm def}}}{{=}}\mathbb{H}^{\sigma}_{0}\times\mathcal{O}\times\mathbb{S}^{2}\times{\mathbb{R}}^{6}.$
@@
(6.3) $\chi(t)=0\ \text{ for }t\leq 1,\qquad\text{ and }\qquad\chi(t)=1\ \text{ for }t\geq 2.$
@@
(6.4) $F^{(r)}(V)\stackrel{{\scriptstyle{\rm def}}}{{=}}\big{[}1-\chi_{3r}(\|v\|_{H^{\sigma}})\big{]}F(V)+\chi_{r}(\|v\|_{H^{\sigma}})H(\mathbf{x},\xi,\eta),$
where, for $V=(v,\mathbf{x},\xi,\eta)\in\mathcal{H}$,
$H(\mathbf{x},\xi,\eta)=(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{4}\,\begin{pmatrix}0\\ \sum_{j=1}^{3}a_{j}\frac{\eta_{j}}{\sqrt{1+|\eta_{j}|^{2}}}\\ \Pi_{\xi}\sum_{j=1}^{3}a_{j}\frac{\eta_{j+3}}{\sqrt{1+|\eta_{j+3}|^{2}}}\\ 0\end{pmatrix},$
$(a_{j})_{j=1}^{3}$ denotes the canonical basis of ${\mathbb{R}}^{3}$ and $\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})$ the distance from $\partial\mathcal{O}$. The exponent used to quantify the distance to the boundary may not be optimal; our choice is guided by the exponent that appears in the main algebraic lemmas below—see Lemma 6.10.
The above truncation (slightly) differs from the one used in [16, Section 6]. Indeed, the additional term $\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})$ ensures that the $\mathbf{x}_{t}$-dynamics remains within $\mathcal{O}$. As in [16], the process $\eta$ is used in combination with the truncation to transfer some ‘randomness’ from the $\eta$-equation to $\mathbf{x}_{t}$ and $\xi_{t}$ via the nonlinearity $\chi_{r}(\|v\|_{H^{\sigma}})H(\mathbf{x},\xi,\eta)$ when $\|v\|_{\mathbb{H}^{\sigma}}$ is large. However, for small values of $\|v\|_{\mathbb{H}^{\sigma}}$, the coupling is not efficient enough, and a hypoellipticity argument has to be exploited; see Subsection 6.5.1.
@@
where $V\in\mathcal{H}$. The global well-posedness of (6.5) follows from standard arguments due to the presence of the truncation, Lemma 5.2 and the smoothness condition $\sigma\geq 2$. For all $r,t\geq 0$ and bounded Borel maps $\phi:\mathcal{H}\to{\mathbb{R}}$, we let
@@
where $V_{t}$ and $V_{t}^{(r)}$ is the solution to (6.2) and (6.5) with initial data $V$, respectively. As in Subsection 3.3, the process $(V_{t})_{t}$ (resp. $(V_{t}^{(r)})_{t}$) is strong Feller in $\mathcal{H}$ if $P_{t}\phi\in C(\mathcal{H};{\mathbb{R}})$ (resp. $P_{t}^{(r)}\phi\in C(\mathcal{H};{\mathbb{R}})$) for all bounded Borel maps $\phi:\mathcal{H}\to{\mathbb{R}}$.
The following result enables us to deduce the strong Feller property of the process $(V_{t})_{t\geq 0}$ from that of the truncated process $(V^{(r)}_{t})_{t\geq 0}$ provided $r$ is sufficiently large.
@@
Fix a bounded Borel observable $\phi:\mathcal{H}\to{\mathbb{R}}$, a time $t\geq 0$ and an initial data $V\in\mathcal{H}$. To prove the claim, we need to show that, for each $\varepsilon>0$, there exists $\delta=\delta(\varepsilon)>0$ such that
@@
Let $V=(v,\mathbf{x},\xi,\eta)\in\mathcal{H}$. By (6.1), it is clear that $V_{t}=(v_{t},\mathbf{x}_{t},\xi_{t},\eta+W^{(\eta)}_{t})$ for $t\geq 0$, where $v_{t}$ is the unique solution to (1.1) and $(\mathbf{x}_{t},\xi_{t})$ the corresponding processes uniquely determined by $v_{t}$ as in (6.1). By Theorem 5.1(1), we can choose $r_{1}\geq r_{0}$ (depending on $\|v\|_{\mathbb{H}^{\sigma}}$) such that, for $\widehat{v}\in\mathbb{H}^{\sigma}_{0}$ satisfying $\|v-\widehat{v}\|_{\mathbb{H}^{\sigma}_{0}}\leq 1$, the solution $\widehat{v}_{t}$ of the stochastic PEs (1.1) with initial data $\widehat{v}\in\mathbb{H}^{\sigma}_{0}$ satisfies
$\mathbf{P}\Big{(}\sup_{s\in[0,t]}\|\widehat{v}_{s}\|_{\mathbb{H}^{\sigma}}\geq r_{1}\Big{)}\leq\frac{\varepsilon}{8\|\phi\|_{L^{\infty}}}.$
@@
we have $\mathbf{P}(\widehat{\tau}<t)\leq\varepsilon/(8\|\phi\|_{L^{\infty}})$. By uniqueness of solutions to the stochastic PEs (cf., Definition 2.1), it follows that $v^{(r)}=v$ on $[0,\tau)\times\Omega$. Hence, for all $r\geq r_{1}$ and $\widehat{V}=(\widehat{v},\mathbf{x},\xi,\eta)\in\mathcal{H}$ such that $\|v-\widehat{v}\|_{\mathbb{H}^{\sigma}}\leq 1$,
@@
Thus, for all $r\geq r_{1}$ and $V^{\prime}\in\mathcal{H}$ satisfying $\|V-V^{\prime}\|_{\mathcal{H}}\leq 1$,
@@
In light of Lemma 6.1, it is enough to prove the Markov process $(V^{(r)}_{t})_{t}$ is strong Feller in $\mathcal{H}$ for $r$ sufficiently large. The latter fact is a consequence of the following result. In its formulation, we use the following notation:
@@
$|DP^{(r)}_{t}\phi(V)h|\lesssim t^{-a_{0}}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{-b_{0}}(1+|\eta|^{b_{0}})\|\phi\|_{L^{\infty}(\mathcal{H})}\|h\|_{T_{V}\mathcal{H}}$
@@
As in [14, Subsection 6.1], the above readily implies the strong Feller property of $(V^{(r)}_{t})_{t}$ with $r\geq r_{0}$. Indeed, fix $\phi\in C^{2}(\mathcal{H};{\mathbb{R}})$ and let $V=(v,\mathbf{x},\xi,\eta),V^{\prime}=(v^{\prime},\mathbf{x}^{\prime},\xi^{\prime},\eta^{\prime})\in\mathcal{H}$ be such that $\|V-V^{\prime}\|_{\mathcal{H}}\leq\varepsilon$ with $\varepsilon\in(0,\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})/2)$. In particular, $\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}^{\prime})\leq\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x})/2$. Let $[0,1]\ni\lambda\to\xi^{\lambda}$ be a geodesic on $\mathbb{S}^{2}$ such that $\xi^{0}=\xi$ and $\xi^{1}=\xi^{\prime}$, and $|\frac{\mathrm{d}}{\mathrm{d}t}\xi^{\lambda}|\lesssim|\xi-\xi^{\prime}|$. Letting $V^{\lambda}=((1-\lambda)v+\lambda v^{\prime},(1-\lambda)\mathbf{x}+\lambda\mathbf{x}^{\prime},\xi^{\lambda},(1-\lambda)\eta+\lambda\eta)$, from Proposition 6.2, for all $t\leq T_{0}$, we have
$|P^{(r)}_{t}\phi(V)-P^{(r)}_{t}\phi(V^{\prime})|$ $\textstyle\leq\int_{0}^{1}\big{|}DP^{(r)}_{t}\phi(V_{\lambda})\frac{\mathrm{d}}{\mathrm{d}\lambda}V^{\lambda}\big{|}\,\mathrm{d}\lambda$
$\textstyle\leq t^{-a_{0}}(\mathrm{dist}_{\partial\mathcal{O}}(\mathbf{x}))^{-b_{0}}(1+|\eta|^{b_{0}})\|\phi\|_{L^{\infty}(\mathcal{H})}\|V-V^{\prime}\|_{\mathcal{H}}.$
Since the right-hand side of the above depends only on $\|\phi\|_{L^{\infty}}$, the previous extends from $\phi\in C^{2}(\mathcal{H};{\mathbb{R}})$ to $\phi$ being only bounded and measurable (see [25, Lemma 7.1.5] for details). The case $t>T_{0}$ follows from the semigroup property and the Feller property (see Proposition 3.1). Hence, $(V^{(r)}_{t})_{t}$ is strong Feller for all $r\geq r_{0}$.
In the rest of this section, we focus on proving Proposition 6.2. Our proof of Proposition 6.2 follows the line of [16, Proposition 6.1] done in Section 6 of the same work. However, adjustments are needed for the stochastic PEs (1.1). In particular, because of the noise degeneracy of the third component $w_{t}$ of the vector field $u_{t}$ (see (2.1)), which creates additional difficulties in the study of the invertibility of the partial Malliavin matrix, see Subsection 6.5.1.
@@ 6.2.1. Basics on Malliavin calculus
In this subsection, we introduce the key concepts in the context of Malliavin calculus, primarily to fix the notation for the subsequent subsections. For a more detailed discussion, the reader is referred to the monograph [62], and for extensions to the Banach-valued case, see [66]. Malliavin calculus is typically introduced in the context of isonormal Gaussian processes [62, Definition 1.1.1]. For our purposes, we focus here on $\mathscr{U}$-cylindrical Brownian motion $\mathscr{W}$ on a separable Hilbert space; see e.g., [8, Definition 2.11] for the definition. Recall that such a cylindrical process is equivalent to an isonormal Gaussian process on $L^{2}(0,T;\mathscr{U})$. To see the correspondence between the latter situation and the one analysed here, let us mention that the sequence of standard independent Brownian motions $(W^{(\mathbf{k},\ell)})_{(\mathbf{k},\ell)\in\mathcal{Z}}$ in (2.12) uniquely induced a $\mathbb{L}^{2}_{0}$-cylindrical Brownian motion via the formula: For $f\in\mathbb{L}^{2}_{0}$ and $0\leq s<t\leq T$,
@@
Using the previous identification, the Malliavin derivative $\mathscr{D}$ can be defined as in [62, Subsection 1.2] or [66, Subsection 2.2] on an Hilbert space $\mathscr{H}$ as an unbounded linear operator
@@
In general, computing the Malliavin derivative is not straightforward. Below, we will need to compute the Malliavin derivative of $V_{T}$, where $V_{T}$ is the solution to the truncated SPDE (6.2) for a fixed $r>0$. However, as shown in [62, Subsection 1.2.1], when $\Omega$ is the standard space of paths, the Malliavin derivative of the random variable $V_{T}$ be computed using the Fréchet derivative of the Itô-map (if available) associated to (6.2), i.e., the map which associated to the noise realisation $W$ the solution $V_{T}(W)$, see [62, Subsection 1.2]. To make this concrete, we need to introduce some more notation. To this end, as in Subsection 3.3 and without loss of generality, we may assume that
@@
Using (6.13), the Fréchet differentiability of the Itô map can be demonstrated as in [41, Proposition 4.1]. For the equality in (6.14), the proof can be obtained by arguing as in [62, Subsection 1.2.1].